0 0
Add a comment Improve this question Transcribed image text
Answer #1
Concepts and reason

A compound is given for which 1HNMR^1{\rm{H - NMR}} with the correct splitting pattern of the proton needs to be given. Equivalent protons are present in chemically same environment, thus, identify the protons which are having similar chemical environment.

Similarly, the number of signals obtained in NMR is also depend on the different type of protons present in the compound. Thee splitting pattern is given by (n+1)\left( {{\rm{n}} + {\rm{1}}} \right) rule and intensity of the peak is shown by the pascals triangle.

Fundamentals

• NMR stands for nuclear magnetic resonance which is a useful spectroscopic technique for determining the type of protons and the number of protons in the compounds.

• A set of protons which are present in same chemical environment, that is, present between the same group of atoms are said to be chemically equivalent protons. All chemically equivalent proton generates 1 signal or 1 peak in 1HNMR^1{\rm{H - NMR}} .

• A set of protons which are present in different chemical environment, that is, present between the different group of atoms are said to be chemically non-equivalent protons. All chemically non-equivalent proton generates distinct signal in 1HNMR^1{\rm{H - NMR}} .

• Splitting of peak: It is also referred as spin-spin splitting. It follows (n+1)\left( {{\rm{n}} + {\rm{1}}} \right) rule, where n is the number of adjacent protons.

Calculate the number of chemically different proton and write the number of signals shown by the given compound in 1HNMR^1{\rm{H - NMR}} as marked in the following structure:

As there are two different sets of protons are present, hence, the number of signals obtained in 1HNMR^1{\rm{H - NMR}} is two.

Show the splitting pattern for the given set of protons following (n+1)\left( {{\rm{n}} + {\rm{1}}} \right) rule as shown below:

a quartet (9)
HH
C CH
“
triplet (t)

Draw the NMR of the of the given compound with the correct representation of its chemical shift values as follows:

NMR H Estimation
3.42
3.42
3Hb
2H
PPM

Ans:

NMR H Estimation
3.42
3.42
3Hb
2H
PPM